OpenStudy (anonymous):

This may be a stupid question...but is 3^2/5^2 the same as 1/(5/3)^2

OpenStudy (ikram002p):

lets see :) 1/(5/3)^2 =1/(5^/3^2)=1*(3^/5^2)=(3^/5^2) any doubt ?

ganeshie8 (ganeshie8):

Ofcourse that works, I think it would be more intereting to ponder on why it works

OpenStudy (anonymous):

oh...thankyou. I forgot that when you divide by a fraction you multiply by the reciprocal

ganeshie8 (ganeshie8):

why \(a/b\) equals \(1/(b/a)\) ?

OpenStudy (anonymous):

Thank you for writing out the steps so I could see it...I think I just had a mind blank and couldn't see that you had to multiply by the reciprocal and so the fraction flips...